# Elasticity without a reference state: continuum mechanics of active tension nets

### v2

@@ Results
To complete the model for plasticity in the continuum, we need to address how cell rearrangements impact the tension metric $\mathbf{g}$. A priori, the change of the adjacency metric $\mathbf{a}$ corresponds to a reparametrization of $\mathbf{g}$ – the tension manifold remains invariant under cell rearrangements (the macroscopic stress depends only on $\mathbf{g}$ and the physical BCs, also remains invariant). Edge flips occur in the triangulation pass between different triangulations that belong to the same stress state. However, T1s are not only an abstract “remeshing” of the tension manifold – they create a new cell interface whose tension dynamics are controlled by specific (bio-)physical processes at the microscopic level.
More precisely, the edge flips kinematically change the tension anisotropy. This is geometrically evident in the discrete setting (as illustrated in Fig. 5 in App. F). In the continuum, where tension anisotropy is measured by $\mu_{\zeta|z}$, the chain rule leads to a contribution $\tfrac{d}{dt}\mu_{\zeta|z}\sim-\partial_{t}\mu_{a}$ in Eq. (.4.2). The T1-induced change of tension anisotropy elicits subsequent dynamics of $\mathbf{g}$ through the processes that govern the microscopic tensions.
Consider the case of a conventional fluid foam as the paradigmatic case for passive T1s. External forces build up macroscopic stress $\sigma$ through isogonal strain while microscopic tensions remain constant. The isogonal strain will cause T1s, which kinematically build up tension anisotropy that cannot be sustained since surface tension is fixed. Instead, the fluid in the interfaces rapidly redistributes, leading to transient viscous stresses and relaxation back to $\mu_{\zeta|z}=0$ concomitant with a relaxation of macroscopic stress $\sigma$. Passive T1s therefore reproduce Maxwell stress relaxation. In the limit where the microscopic timescale on which viscous stresses relax is zero, $\mathbf{g}\equiv\mathbf{a}$ and $\zeta=z$ in a conventional foam. Therefore,
$\partial_{t}\mu_{g}=\partial_{t}\mu_{a}=\gamma_{\mathrm{T1}}\mu_{z|w}\approx\gamma_{\mathrm{T1}}\frac{\sigma_{ww}}{2p_{0}}$ (56)
since deviatoric strain $\mu_{z|w}$ determines the stress via Eq. (.2.4) (of the the linearized theory). On the other hand, tension metric dynamics determines the change in stress. Indeed, for fixed physical positions $w$, the Beltrami equation (10) implies $\frac{\delta\mu_{z|w}}{\delta\mu_{g}}=-1$, and therefore, $\partial_{t}\sigma_{ww}=-\partial_{t}\mu_{g}/(2p_{0})$. Together, these two effects result in effective Maxwell relaxation $\partial_{t}\sigma\sim-\sigma$. (While T1s relax only the anisotropic stress, due to mechanical balance Eq. (22), this also removes isotropic stress gradients.) For the reference-state velocity field in the purely passive case, stress relaxation and the “adiabatic flow” Eq. (43) lead to an effective Stokes flow:
@@
$\frac{p_{0}}{2}\Delta\dot{z}_{\beta}=\gamma_{\mathrm{T1}}\partial_{\alpha}\!\left(\sigma^{\mathrm{tot}}_{\alpha\beta}-\tfrac{1}{2}\sigma^{\mathrm{tot}}_{\gamma\gamma}\delta_{\alpha\beta}\right)$ (57)
The drive results from the anisotropic macroscopic stress, and the T1 rate sets the relaxation rate. Going from $\dot{z}_{\beta}$ to the physical flow $\dot{w}_{\beta}$ requires additionally solving Eq. (44). If $\dot{K}_{g}=0$, one simply has $\dot{w}_{\beta}=\dot{z}_{\beta}$.
Let us next consider the purely active case, $w=z$, i.e. $\mu_{z|w}=0$. Here, T1s happen when $|\mu_{\zeta|z}|$ increases either by positive feedback [29, 16] or by an “external field” like a morphogen gradient [31, 30]. Substituting Eq. (.4.3) into Eq. (.4.2) reveals that T1s kinematically relax tension anisotropy: $\tfrac{d}{dt}\mu_{\zeta|z}=\partial_{t}\mu_{\zeta|z}-\gamma_{\mathrm{T1}}\mu_{\zeta|z}$ (suppressing the advection and co-rotation terms). Steady active plastic flow results when driving and relaxation balance at a finite T1 rate. (Appendix G.3 describes a concrete model where tensions are governed by positive mechanical feedback [30].) Active flow can take place at vanishing macroscopic stress and is therefore fundamentally different from conventional Maxwell viscoelasticity. Instead, it is characterized by sustained microscopic tension anisotropy, experimentally detectable as anisotropic junctional myosin [32].
We conclude that both active and passive T1s relax their “driving anisotropy”: tension anisotropy in the former case and macroscopic stress anisotropy in the latter. We can combine both in a unified phenomenological model for the T1-contribution to the dynamics of $\mathbf{g}$
@@
$\left(\partial_{t}\mu_{g}(\zeta,\bar{\zeta})\right)_{\mathrm{T1}}=\gamma_{\mathrm{T1}}e^{-i2\phi_{\zeta|z}}\mu_{z|w}=-\gamma_{\mathrm{T1}}\frac{\delta|\mu_{z|w}|^{2}}{\delta\bar{\mu}_{g}},$ (58)
Where the second equality follows from Eq. (.4.3). Additional contributions can arise from mechanical feedback and “external fields” driving $\mu_{g}$, while in a passive tissue (or conventional foam), only T1s contribute.
@@
To summarize, the present framework describes a tissue by two metrics $\mathbf{a},\mathbf{g}$ that can be controlled independently. A kinematical prescription determines what happens during topological change (dynamics of $\mathbf{a}(t)$, keeping $\mathbf{g}(t)$ fixed). The dynamics of the tension metric $\mathbf{g}$ depend on the specific model or biological context of interest. We showed that minimal and microscopically motivated choice, Eq. (58), naturally gives rise to Maxwell-like stress relaxation through passive T1s while active T1s relax tension anisotropy.
@@ Discussion
We have presented a continuum theory for active tension nets, a paradigmatic example of “active elasticity”, where the local stress, rather than the local target shape, is the physical (or biological) control parameter. We showed how the intrinsic configuration of active stress controls the large-scale shape of a tissue. Adiabatic changes of stress drive active plastic flow, described as a QC flow governed by two coupled Beltrami equations, Eqs. (40) and (52).
@@
Crucially, many microscopic tension triangulations specify the same tension metric. They correspond to reparametrizations of the tension manifold and do not affect physical stress. Put differently, the macroscopic (i.e., tissue-scale) mechanical state only depends on the shape of the tension manifold, not how it is discretized by “painting” individual tension triangles onto it, like in a finite-element discretization of a partial differential equation. In particular, different triangulations of a region in the plane all define the same tension metric – only the shape of the region’s boundary is relevant. In that sense, for a flat tension metric, the continuum theory is “holographic”: the configuration of the tissue is encoded in the shape of the boundary and the traction forces acting there. This holographic nature of emergent elasticity is also relevant in granular matter, where boundary forces constrain the ensemble of self-stresses in the bulk [33].
@@
Of course, the microscopic cell shapes depend on the tension triangulation. Moreover, tensions are ultimately sensed and controlled at the cellular level. Therefore, their dynamics depend on the microscopic tension configuration. A continuum model, therefore, requires information in addition to $\mathbf{g}$. The most important piece is the adjacency metric $\mathbf{a}$, whose continuum dynamics captures cell rearrangements. While the tension metric $\mathbf{g}$ determines the macroscopic mechanics, the local tension anisotropy is encoded in the “difference” of $\mathbf{g}$ and $\mathbf{a}$ – the shape of tension triangles. Overall, the present theory comprises the minimal description of tension nets in the continuum: The tension metric $\mathbf{g}({\bm{\xi}})$ (equivalently, isothermal embedding $z(\xi,\bar{\xi})$ and curvature $K_{g}$) and the adjacent metric $\mathbf{a}({\bm{\xi}})$, for a total of six degrees of freedom. The systematic description of stress and network topology in terms of Riemannian geometry and QC maps may prove useful in the study of other 2D materials with dynamic microscopic structure, such as granular matter.
Active plastic flow.
@@
To describe tissue flow driven by adiabatic dynamics of active stress, one tracks the reference state $z(\xi,\bar{\xi},t)$ of the evolving tension metric $\mathbf{g}(t)$. The time evolution $\dot{z}$ is governed by a Beltrami flow, Eq. (40). Large deformations entail topological cell rearrangement. This raises the challenge of incorporating discrete topology into a continuum theory. On a cell level, topological information can be “geometrized” via the so-called circle packing of the cell adjacency graph, whose existence and uniqueness is guaranteed by the Koebe–Andreev–Thurston theorem [34]. As explained in the companion paper [15], the circle packing is a “discrete” conformal embedding of the equilateral triangulation formed by the adjacency graph. In the continuum, it becomes the isothermal embedding $\zeta(\xi,\bar{\xi})$ of the adjacency metric $\mathbf{a}$. Tension dynamics, plastic flow, and elastic deformation can thus be formulated as two coupled Beltrami flows, Eqs. (40), and (52). This principled approach naturally incorporates the local tension configuration via the Jacobian $\partial_{\bm{\zeta}}\mathbf{z}$.
In conventional models of tissue flow, imbalanced active stresses act against viscous dissipation [32, 35, 36]. In the present theory, by contrast, it is the adiabatic change of active stress that drives flow. This has implications for connecting cell-level models to their large-scale behavior. In this picture, a steady-state tension anisotropy emerges from the balance of anisotropy-increasing processes (such as mechanical feedback or morphogen “drive”) and relaxation by T1s. By contrast, the macroscopic stress remains isotropic (in the absence of external forces). Tension anisotropy has previously been identified with an anisotropic active stress tensor [32]. These discrepant predictions could be tested by laser ablations on a supracellular scale.
@@
In the “active plasticity” model, the tissue remodels from one solid configuration to the next – it flows while remaining solid, able to resist external shear forces [^7] Note, however, that the yield strain becomes highly anisotropic during active plastic flow. Due to the ongoing T1 transitions, there is a large number of cells near the T1 threshold, and the yield strain parallel to tension anisotropy can become very small.. By contrast, in a conventional active fluid, flow results from viscous dissipation, and the system is instantaneously fluid and cannot resist shear forces.
@@
On a formal level, the present theory has parallels with metric elasticity, where a Riemannian reference metric represents the rest lengths between adjacent material points [38], and deviations from these rest lengths in the physical configuration determine the elastic stress via a constitutive law. By contrast, the tension metric specifies the active stress (tension flux) between adjacent points. A stress-free reference state and an effective stress-strain relationship then emerge from the intrinsic stress configuration. The physical stress tensor is determined by the embedding of the tension metric given by the cell positions, i.e., the arrangement of active force dipoles in physical space. The present theory, therefore, furnishes the continuum limit of cell-level models dominated by active mechanics. By contrast, conventional metric elasticity is well-suited to describe the continuum limit of passive models (generalized spring-networks). For instance, in the context of the well-known “area-perimeter” vertex model [7], Ref. [8] defined a network tensor that plays the role of the elastic reference metric. Plastic flow can be modeled by relaxational dynamics of the reference metric [39].
An attractive feature of metric elasticity is its variational formulation. Our theory also admits a variational form: The isothermal embedding $\mathbf{z}({\bm{\xi}})$ of the tension metric can be characterized by a pseudo-energy, formally similar to metric elasticity. Since $\mathbf{z}({\bm{\xi}})$ is a conformal map from the tension manifold into the Euclidean plane, its components must be harmonic functions. Equivalently, $z(\xi,\bar{\xi})$ must be a minimum of the Dirichlet pseudo-energy $E_{D}=\int d^{2}\xi\sqrt{g}\ g^{\alpha\beta}\partial_{\alpha}{\bar{z}}(\xi)\partial_{\beta}{z}(\xi)$ [40] (with suitable BCs, see App. C). However, $E_{D}$ derives from the geometry of force balance, not from a material law, and does not determine the stress tensor $\sigma_{\alpha\beta}$. A better understanding of the variational formulation of active mechanics is an interesting avenue for future work.
Tissues on curved and dynamic surfaces.
For simplicity, the present manuscript considers a tissue embedded in the plane. However, the geometric nature of the theory implies that it easily generalizes to tissues on curved surfaces, such as the ellipsoidal blastoderm of the Drosophila. To this end, the Euclidean metric $\delta_{\alpha\beta}$ in Eq. (11) must be replaced by the metric of the curved physical space. This is most conveniently done by parameterizing physical space with (yet another set of) isothermal coordinates. The reference configuration $z(\xi,\bar{\xi})$ is found as a conformal map between two curved surfaces – the tension- and physical manifold – and the mechanics is controlled by their relative geometry – e.g., the difference of Gaussian curvatures in Eq. (12).
@@
Gaussian curvature of the tension metric leads to pressure differentials, similar to residual stresses in metric elasticity due to an incompatible target metric. Interestingly, many tissues can avoid large pressure differentials and the potential instabilities they entail, like cell extrusion or rupture. On the cell level, pressures can be locally controlled via (negative) feedback loops  [^8] The pressure feedback discussed here operates quasi-statically and requires a non-zero bulk modulus to build up pressure gradients. By contrast, in a compressible tissue, tension incompatibility causes secular elongation or contraction of cell-cell interfaces, requiring a different stabilizing mechanism (see Ref. [11]). coupling tensions and intracellular pressure. A minimal mathematical model for such feedback is $\partial_{t}\mathbf{g}=-\beta(\Delta\log p/p_{0})\mathbf{g}=-\beta K_{g}\,\mathbf{g}$, where $\beta$ is a feedback coefficient and we used Eq. (12). Cells can hence implement a “flattening” Ricci flow [43] by modulating their overall contractile activity in response to intracellular pressure. In contrast to, e.g., Frank elasticity in liquid crystals, Ricci flow only acts on the “curvature part” of $\mathbf{g}$ – the tension manifold remains free to remodel in-plane to drive morphogenetic flows.
On the tissue level, pressure can be controlled via feedback through cell proliferation and growth [44, 1, 45], which in our formulation would act on the target cell density $n({\bm{\xi}})$ (Eq. (17)). Analyzing such “adaptive” feedback loops, and the combination of robustness and plasticity they confer, is an interesting route for future research.
More generally, our findings have interesting consequences for how cells can sense and react to mechanical cues. We showed that the macroscopic stress $\sigma$ is independent of the local tension configuration, and, therefore, “invisible” on a cell level. Instead, the cell-level “observables” are the tension anisotropy $\mu_{\zeta|z}$ and the intracellular pressure $p$. Cell shape anisotropy could be sensed by intracellular structures such as the nucleus, intermediate filaments, or microtubules.
@@
The link between the continuum theory and cell level (the discrete ATN model) is discussed in detail in a companion paper [15]. There, we exploit a theory of discrete conformal maps [46, 47] to bridge the gap between the continuum theory of the present manuscript and the mechanics at the cellular level. This theory constructs a discrete equivalent of the isothermal embedding $z(\xi,\bar{\xi})$. The isogonal (curl-free) and conformal modes identified by the continuum analysis reappear as the geometric soft modes which respect the microscopic force balance constraints [48, 11, 49], while the Liouville equation for the curvature of the tension manifold leads to a generalized von Neumann law for the cell pressures. Coarse-graining the microscopic stress confirms the stress-transformation law Eq. (21). Thus, the “top-down” analysis presented here is independently confirmed by “bottom-up” coarse-graining.
Similarly, (active) plasticity, mediated by cell rearrangements, is captured by remodeling of the cell-adjacency graph. In the continuum, this graph is coarse-grained to the adjacency metric $\mathbf{a}$. The circle packing provides a discrete equivalent to the $\mathbf{a}$-isothermal coordinates (see paragraph on “active plastic flow” above). In the discrete setting, the condition for T1 transitions can be calculated explicitly from purely geometric considerations (a generalized “Delaunay criterion”), justifying the phenomenological T1 rate $\propto\mu_{\zeta|w}$ in amorphous tissue with negligible micro-structural order (Eq. (58)). Micro-structural order can strongly influence T1 transitions, in which case additional order parameters become relevant [16]. A continuum theory for these order parameters and their relation to topological defects in the microscopic topology is an interesting direction for future research.

### v3

@@ Results
To complete the model for plasticity in the continuum, we need to address how cell rearrangements impact the tension metric $\mathbf{g}$. A priori, the flip flow due to cell rearrangements is a reparametrization– the tension manifold remains invariant (the macroscopic stress, depending only on $\mathbf{g}$ and the physical BCs, also remains invariant). Edge flips in the triangulation pass between different triangulations that belong to the same stress state. However, T1s are not only an abstract “remeshing” of the tension manifold; they create a new cell interface whose tension dynamics are governed by specific (bio-)physical processes at the microscopic level. More precisely, by Eq. (.4.2), edge flips change the tension anisotropy $\mu_{\zeta|z}$ (this is geometrically evident in the discrete setting; Fig. 8). The T1-induced change of tension anisotropy can then elicit subsequent dynamics of $\mathbf{g}$ through the processes that govern the microscopic tensions.
For simplicity, in this section we linearize, assuming small strain and tension anisotropy, $|\zeta-z|,|z-w|\ll 1$, and further $\lambda_{g}\approx 1$. We can thus neglect the distinctions between the different coordinate systems $\zeta,z,w$, and express everything $z$-coordinates. As a minimal model, we propose that passive T1-transitions remodel the tension metric, in addition to the active “drive” provided by $\mathbf{m}$.
$\partial_{t}\log\mathbf{g}=\mathbf{m}+\mathbf{R}_{\mathrm{pass}}$ (57)
Hence, for the Beltrami coefficient, $\partial_{t}\mu_{g}=m_{zz}/4+R_{\mathrm{pass}}$.
As we now show, Eq. (57) captures the physically expected behavior for both active and passive plastic flow. Using the passive T1-rate Eq. (56b), and relating deviatoric strain $\mu_{z|w}$ to the stress via Eq. (20):
$\partial_{t}\mu_{g}=\frac{m_{zz}}{4}+\gamma_{\mathrm{T1}}\mu_{z|w}=\frac{m_{zz}}{4}+\gamma_{\mathrm{T1}}\frac{\sigma_{zz}}{2p_{0}\lambda_{g}}$ (58)
By a slight abuse of notation, we write $\sigma_{zz}$ for the elastic stress in $z$-coordinates, which, of course, depends on the deformation $w$. On the other hand, changing the tension metric at fixed physical configurations changes the stress. We evaluate the Cauchy stress Eq. (4) in the current configuration, but using the time-incremented tension metric $\mathbf{g}(t+dt)$:
$\bm{\sigma}=p_{0}\epsilon^{T}\!\cdot\!\sqrt{\mathbf{g}(z)}\cdot\epsilon=p_{0}\epsilon^{T}\!\cdot\!\sqrt{\mathbb{I}+(\partial_{t}\mathbf{g})dt}\cdot\epsilon$ (59)
(Strictly speaking, one should expand around $\mathbf{g}(w)$ and not $\mathbf{g}(z)$; we neglect the difference.) To linear order:
$\partial_{t}\sigma_{zz}\big|_{w=\mathrm{const.}}=-\frac{p_{0}\lambda_{g}}{2}(\partial_{t}\mu_{g})$ (60)
The minus sign comes from the $\pi/2$ rotation $\epsilon$. Together, Eqs. (58), (60) result in effective Maxwell relaxation, with a source provided by $\mathbf{m}$:
$\partial_{t}\sigma_{zz}\big|_{w=\mathrm{const.}}=-\gamma_{\mathrm{T1}}\sigma_{ww}-2p_{0}m_{zz}$ (61)
Eq. (61) describes how passive T1s relax stress when the cell positions $w$ are pinned. Note that on the microscopic level, this corresponds to pinning cell centroids $\mathbf{r}_{i}$; cell vertices $\mathbf{r}_{ijk}$ are still free to move to establish local force balance. While T1s relax only the anisotropic stress, due to mechanical balance Eq. (22a), vanishing anisotropic stress also implies uniform isotropic stress gradients.
For the reference-state velocity field, the combination of stress relaxation and active driving manifests as an effective Stokes flow [cf. Eq. (43)]:
@@
$\Delta\dot{z}_{\beta}$ $=\frac{2\gamma_{\mathrm{T1}}}{p_{0}}\partial_{\alpha}\!\left(\sigma_{\alpha\beta}-\tfrac{1}{2}\sigma_{\gamma\gamma}\delta_{\alpha\beta}\right)$
@@
$\quad+\partial_{\alpha}\left(m_{\alpha\beta}-\tfrac{1}{2}m_{\gamma\gamma}\delta_{\alpha\beta}\right)$ (62)
The anisotropic macroscopic stress enters with an effective viscosity $p_{0}/(2\gamma_{\mathrm{T1}})$. Note that in many cases, $\gamma_{\mathrm{T1}}$ may vanish below a critical yield strain [38]. Eq. (62) thus combines active flow due to tension dynamics $\mathbf{m}$ with viscous relaxation of elastic stress due to passive T1s. In the limiting cases discussed so far (free boundaries with flat tension metric; pinned cells) are particularly simple since one does not need to solve the effective elastic problem Eq. (44) for the physical flow $\dot{w}_{\beta}$ (Sect. .3.3). In general, one must solve this non-local problem, as formally stated in Eq. (39), find the physical configuration $w$ and thus the T1 rate $R=\gamma_{\mathrm{T1}}\mu_{\zeta|w}$.
We now explore the physical consequences of the proposed model in the purely active and passive cases. Consider a conventional fluid foam, where $\mathbf{m}=0$, as the paradigmatic case for passive T1s. External forces can build up deviatoric strain and thus macroscopic stress $\bm{\sigma}$, while microscopic tensions remain constant. Eventually, strain causes T1s that relax macroscopic stress. Eq. (62) reproduces this physical intuition. Further, since all tensions equal the fluid’s surface tension, $\mathbf{g}=\mathbf{a}$: T1s should directly modify the tension metric. This is precisely ensured by the dynamics Eq. (57).
Let us next consider the purely active case, $w=z$, i.e. $\mu_{z|w}=0$. Here, T1s happen when $|\mu_{\zeta|z}|$ increases due to driving by $m_{zz}$ (for example, positive mechanical feedback [38, 19] or a morphogen gradient [40, 39]). Substituting Eq. (57) into Eq. (.4.2) shows that T1s relax tension anisotropy:
$\frac{d}{dt}\mu_{\zeta|z}=\tfrac{1}{4}m_{zz}-R_{\mathrm{act}}$ (63)
Active T1s thus limit the tension anisotropy a tissue can attain. Steady active plastic flow results when driving and relaxation balance at a finite T1 rate. Active flow can take place at vanishing macroscopic stress and is therefore fundamentally different from conventional Maxwell viscoelasticity. Instead, it is characterized by sustained microscopic tension anisotropy, experimentally detectable as anisotropic junctional myosin [15].
@@
To summarize, the present framework describes a tissue by two independently controlled metrics $\mathbf{a},\,\mathbf{g}$. A kinematical prescription determines what happens during topological change (dynamics of $\mathbf{a}(t)$, keeping $\mathbf{g}(t)$ fixed). The dynamics of the tension metric $\mathbf{g}$ depend on the specific model or biological context of interest. We showed that minimal and microscopically motivated choice, Eq. (57), naturally gives rise to Maxwell-like stress relaxation through passive T1s while active T1s relax tension anisotropy.
@@ Discussion
We have presented a continuum theory for active tension nets, a paradigmatic example of an active solid, where the local stress, rather than the local target shape, is the physical (or biological) control parameter. Microscopic tensions are purely active because passive stresses rapidly dissipate through molecular turnover. This distinguishes the type of active solid studied here from previous work where active stresses are added on top of a conventional (passive) elastic background [41, 42, 43] (see below).
@@
Crucially, many microscopic tension triangulations specify the same tension metric. They correspond to reparametrizations of the tension manifold and do not affect physical stress. Put differently, the macroscopic (i.e., tissue-scale) mechanical state only depends on the shape of the tension manifold, not how it is discretized by “painting” individual tension triangles onto it, like in a finite-element discretization of a partial differential equation. In particular, different triangulations of a region in the plane all define the same tension metric – only the shape of the region’s boundary is relevant. For a flat tension metric, the tissue configuration is fully encoded in the boundary shape and the traction forces. This “holographic” nature of emergent elasticity is also relevant in granular matter, where boundary forces constrain the ensemble of self-stresses in the bulk [44].
@@
The microscopic cell shapes depend on the tension triangulation. Moreover, tensions are ultimately sensed and controlled at the cellular level. Therefore, their dynamics depend on the microscopic tension configuration. A continuum model, therefore, requires information in addition to $\mathbf{g}$. The most important piece is the adjacency metric $\mathbf{a}$, whose continuum dynamics captures cell rearrangements. While the tension metric $\mathbf{g}$ determines the macroscopic mechanics, the local tension anisotropy is encoded in the “difference” of $\mathbf{g}$ and $\mathbf{a}$ – the shape of tension triangles. Overall, the present theory comprises the minimal description of tension nets in the continuum: The tension metric $\mathbf{g}({\bm{\xi}})$ (equivalently, isothermal embedding $z(\xi,\bar{\xi})$ and curvature $K_{g}$) and the adjacent metric $\mathbf{a}({\bm{\xi}})$, for a total of six degrees of freedom. The systematic description of stress and network topology in terms of Riemannian geometry and QC maps may prove useful to study other 2d materials with dynamic microscopic structure, such as granular matter.
Active plasticity.
@@
To describe tissue flow driven by adiabatic dynamics of active stress, one tracks the reference state $z(\xi,\bar{\xi},t)$ of the evolving tension metric $\mathbf{g}(t)$. The resulting QC adiabatic flow $\dot{z}$ is governed Eq. (42). Large deformations entail topological cell rearrangement. This raises the challenge of incorporating discrete topology into a continuum theory. On a cell level, topological information can be “geometrized” via the so-called ”circle packing” of the cell adjacency graph, whose existence and uniqueness is guaranteed by the Koebe–Andreev–Thurston theorem [45]. As explained in the companion paper [18], the circle packing is a “discrete” conformal embedding of the equilateral triangulation formed by the adjacency graph. In the continuum, it becomes the isothermal embedding $\zeta(\xi,\bar{\xi})$ of the adjacency metric $\mathbf{a}$, and cell rearrangement becomes the flip flow $\dot{\zeta}$. Tension dynamics, plasticity, and elastic deformation can thus be formulated as two coupled QC flows, Eqs. (42), and (52). This principled approach naturally incorporates the local tension configuration via the Jacobian $\partial_{\bm{\zeta}}\mathbf{z}$.
Role and nature of active stress.
Many continuum models for active and tissue mechanics posit a phenomenological active stress tensor $\bm{\sigma}^{\mathrm{act}}$ [15, 16, 17], assumed to be proportional to the local motor molecule concentration. Flow is driven by the divergence of $\bm{\sigma}^{\mathrm{act}}$, balanced against the passive stress $\bm{\sigma}^{\mathrm{pass}}$ of a visco-elastic “background”. The balance of motor-generated force and viscosity sets the flow speed, and the tissue is instantaneously fluid (zero shear modulus).
By contrast, in the present model, there is no passive shear stress (all viscous and passive contributions are assumed to relax rapidly). Local motor molecule concentration is encoded in the tension metric $\mathbf{g}$. The resulting tensile stress $\bm{\sigma}$ depends both on $\mathbf{g}$ and the embedding $\mathbf{r}({\bm{\xi}})$ of active force dipoles into physical space via the stress-metric relation. It is therefore not consistent to prescribe an active stress a priori, independent of the cell configuration. The companion paper confirms this result by explicitly calculating the coarse graining the stress via the Batchelor formula (App. A gives a “preview”).
@@
Plastic flow is caused by tension metric dynamics, and its speed is set by the rate of change of motor molecule concentration. The flow is adiabatic: the stress remains balanced throughout, and the system remodels from one solid configuration to the next. The tissue thus flows while remaining solid  [^8] Note, however, that the yield strain becomes highly anisotropic during active plastic flow. Due to the ongoing T1 transitions, there is a large number of cells near the T1 threshold, and the yield strain parallel to tension anisotropy can become very small.. Notably, adiabatic flow takes the form of an effective Stokes equation, as in previous models describing tissue flow as viscous Stokes flow.
These distinctions have experimentally verifiable consequences. First, the present model predicts that the macroscopic stress remains isotropic during tissue flow (in the absence of external forces). A steady-state tension anisotropy, distinct from stress anisotropy, can emerge from the balance of anisotropy-increasing processes (such as mechanical feedback or morphogens) and resetting by T1s. This could be tested by laser ablations on supracellular scales. Second, it suggests that the speed of tissue flow is set not by the absolute levels of motor molecules and active tensions, but by their rate of change. These predictions rely on the separation of timescales between morphogenetic flow and relaxation of passive stress and mechanical equilibration. It will be important to explore what happens if this assumption is relaxed.
@@
On a formal level, the present theory has parallels with metric elasticity, where a Riemannian reference metric represents the rest lengths between adjacent material points [47, 48, 49], and deviations from these rest lengths in the physical configuration determine the elastic stress via a constitutive law. By contrast, the tension metric specifies the active stress (tension flux) between adjacent points. A stress-free reference state and an effective stress-strain relationship then emerge from the intrinsic stress configuration. The stress-metric relation determines the physical stress tensor in terms of the embedding of the tension metric given by the cell positions, i.e., the arrangement of active force dipoles in physical space. The present theory, therefore, furnishes the continuum limit of cell-level models dominated by active mechanics. By contrast, conventional metric elasticity is well-suited to describe the continuum limit of passive models (generalized spring-networks). For instance, in the context of the well-known “area-perimeter” vertex model [7], Ref. [50] defined a network tensor that plays the role of the elastic reference metric. Plasticity can be modeled by relaxational dynamics of the reference metric [48].
Tissues on curved, topologically non-trivial, and dynamic surfaces.
For simplicity, the present manuscript considers a tissue patch with a disk topology, embedded in the plane. If the tissue topology is non-trivial, new geometric modes and invariants appear. For example, on an annulus, there is a curl-free vector field that cannot be written as a potential gradient [51], and the annulus’ “aspect ratio” is conformally invariant [25]. Due to the important role of curl-free and conformal modes in our theory, this may have interesting physical consequences to be explored in future work. The geometric nature of the theory implies that it generalizes easily to tissues on curved surfaces, such as the ellipsoidal blastoderm of the Drosophila. To this end, the Euclidean metric $\delta_{\alpha\beta}$ in Eq. (11) must be replaced by the metric of the curved physical space. This is most conveniently done by parameterizing physical space with (yet another set of) isothermal coordinates. The reference configuration $z(\xi,\bar{\xi})$ is found as a conformal map between two curved surfaces – the tension- and physical manifold – and the mechanics is controlled by their relative geometry – e.g., the difference of Gaussian curvatures in Eq. (12).
@@
Gaussian curvature of the tension metric leads to pressure differentials, similar to residual stresses in metric elasticity due to an incompatible target metric. Interestingly, many tissues can avoid large pressure differentials and the potential instabilities they entail, like cell extrusion or rupture. On the cell level, pressures can be locally controlled via (negative) feedback loops  [^9] The pressure feedback discussed here operates quasi-statically and requires a non-zero bulk modulus to build up pressure gradients. By contrast, in a compressible tissue, tension incompatibility causes secular elongation or contraction of cell-cell interfaces, requiring a different stabilizing mechanism (see Ref. [10]). coupling tensions and intracellular pressure. A minimal mathematical model for such feedback is $\partial_{t}\mathbf{g}=-\beta(\Delta\log p/p_{0})\mathbf{g}=-\beta K_{g}\,\mathbf{g}$, where $\beta$ is a feedback coefficient and we used Eq. (12). Cells can hence implement a “flattening” Ricci flow [54] by modulating their overall contractile activity in response to intracellular pressure. In contrast to, e.g., Frank elasticity in liquid crystals, Ricci flow only acts on the “curvature part” of $\mathbf{g}$ – the tension manifold can freely remodel in-plane to drive morphogenetic flows.
On the tissue level, pressure can be controlled via feedback through cell proliferation and growth [55, 1, 56], which in our formulation would act on the target cell density $n({\bm{\xi}})$; Eq. (17). Analyzing such “adaptive” feedback loops, and the combination of robustness and plasticity they confer, is an interesting direction for future research.
More generally, our findings have implications for how cells can sense and react to mechanical cues. We showed that the macroscopic stress $\sigma$ is independent of the local tension configuration, and, therefore, “invisible” on a cell level. Instead, the cell-level “observables” are the tension anisotropy $\mu_{\zeta|z}$ and the intracellular pressure $p$. Cell shape anisotropy could be sensed by intracellular structures such as the nucleus, intermediate filaments, or microtubules.
@@
The link between the continuum theory and cell level (the discrete ATN model) is discussed in detail in a companion paper [18]. There, we exploit a theory of discrete conformal maps [57, 58] to bridge the gap between the continuum theory of the present manuscript and the mechanics at the cellular level. This theory constructs a discrete equivalent of the isothermal embedding $z(\xi,\bar{\xi})$. The isogonal (curl-free) and conformal modes identified by the continuum analysis reappear as the geometric soft modes which respect the microscopic force balance constraints [59, 10, 60], while the Liouville equation for the tension curvature leads to a generalized von Neumann law for the cell pressures. Coarse-graining the microscopic stress confirms the stress-metric relation Eq. (21). Thus, the “top-down” analysis presented here is independently confirmed by “bottom-up” coarse-graining.
Similarly, (active) plasticity, mediated by cell rearrangements, is captured by remodeling of the cell-adjacency graph. In the continuum, this graph is coarse-grained to the adjacency metric $\mathbf{a}$. The circle packing provides a discrete equivalent to the $\mathbf{a}$-isothermal coordinates (see paragraph on “active plastic flow” above). In the discrete setting, the condition for T1 transitions can be calculated explicitly from purely geometric considerations (a generalized “Delaunay criterion”), justifying the phenomenological T1 rate $\propto\mu_{\zeta|w}$ in amorphous tissue with negligible micro-structural order; Eq. (54). Micro-structural order can strongly influence T1 transitions, in which case additional order parameters become relevant [19]. A continuum theory for these order parameters and their relation to topological defects in the microscopic topology is an interesting direction for future research.